Please don't copy another solution. Make it original!

From 1998 to 2008, there were 11,590 NBA regular season games. Home teams won 7,021 games while losing 4,569 games.

a. Use these statistics to construct a 95% confidence interval for the proportion of NBA games won by the home team.

​​​​​​​b. (.59688 , .61468)

I am 95% confident that the true confidence interval is between (.59688 , .61468).

c. ​​​​​​​Technical question: Is this an independent data set?

d.​​​​​​​ Should we trust our margin of error here?
Answer #1

a.
TRADITIONAL METHOD
given that,
possible chances (x)=7021
sample size(n)=11590
success rate ( p )= x/n = 0.60578
I.
sample proportion = 0.60578
standard error = Sqrt ( (0.60578*0.39422) /11590) )
= 0.00454
II.
margin of error = Z a/2 * (standard error)
where,
Za/2 = Z-table value
level of significance, α = 0.05
from standard normal table, two tailed z α/2 =1.96
margin of error = 1.96 * 0.00454
= 0.0089
III.
CI = [ p ± margin of error ]
confidence interval = [0.60578 ± 0.0089]
= [ 0.59688 , 0.61468]
-----------------------------------------------------------------------------------------------
DIRECT METHOD
given that,
possible chances (x)=7021
sample size(n)=11590
success rate ( p )= x/n = 0.60578
CI = confidence interval
confidence interval = [ 0.60578 ± 1.96 * Sqrt ( (0.60578*0.39422) /11590) ) ]
= [0.60578 - 1.96 * Sqrt ( (0.60578*0.39422) /11590) , 0.60578 + 1.96 * Sqrt ( (0.60578*0.39422) /11590) ]
= [0.59688 , 0.61468]
-----------------------------------------------------------------------------------------------
interpretations:
1. We are 95% sure that the interval [ 0.59688 , 0.61468] contains the true population proportion
2. If a large number of samples are collected, and a confidence interval is created
for each sample, 95% of these intervals will contains the true population proportion
b.
95% confidence interval for the proportion of NBA games won by the home team is [ 0.59688 , 0.61468]
c.
yes,
an independent data set because it is random sample
d.
margin of error = 0.0089